what two whole numbers when added together the sum is 123,when i subtract the lesser number from the greater number the difference is 45 ?
Answer by josmiceli(19441)

A and B are the whole numbers
A + B = 123
A - B = 45
-A = -A subtract a from both sides
-B = 45 -A
B = A -45 multiply both sides by -1
A + (A-45) = 123 substitution
2A -45 = 123 add 45 to both sides
2A = 168 divided both sides by 2
A = 84
84 - B = 45 substitution
-B = -39 subtract 84 from both sides
B = 39 multiply both sides by -1
A= 84, B=39
84 + 39 = 123 check answers
84 - 39 = 45